Status Report

NOAA GEOS SATOPS Morning Report 9 Jan 2001

By SpaceRef Editor
January 9, 2001
Filed under ,

SATOPS MORNING REPORT – January 9, 2001

POES

No change in the status of the POES spacecraft.

NOAA-15 rev 13815 / W at 0006Z on January 9:  Spacecraft commanding
to place MIRP in AVHRR sync w/rephase enabled and then to place MIRP back
in AVHRR sync w/rephase disabled.

POES operations were nominal over the past 24 hours.

DMSP

OPS45 rev 02818 / FAIR at 1710Z on January 8:  Spacecraft commanding
to turn off the SSM/T-1 and SSM/T-2 (sounder/profiler package) with engineering’s
direction due high temperature limits exceeded.  This action will
prevent the instruments from incurring permanent damage.

DMSP operations were nominal over the past 24 hours.

GOES

No change in the status of the GOES spacecraft.

At 1113Z on January 9:  GOES-10 East/West maneuver performed nominally.

GOES operations were nominal over the past 24 hours.

Other:  Total lunar eclipse at 1950Z today (2:50pm local) and Wallops
power system breaker testing occuring this morning.

 

Questions or comments to [email protected]

SpaceRef staff editor.